### v1

@@ 2.2 Architectures of GNN-agglomeration model for heterogeneous domains
In this section, we further extend the proposed approach to perform mesh agglomeration on heterogeneous domains with different physical parameters. Specifically, this model can be exploited in different situations such as:
  • Heterogeneous domains, where the main goal of the algorithm is to maintain as much as possible the distinction between the different regions still ensuring an acceptable shape for the agglomerated elements.
  • Domains with micro-structures, which is a more complex situation because there is not a precise criterion to perform the agglomeration, and the algorithm aims to understand how to group the tetrahedra to obtain a better grid for numerical methods.
The state-of-the-art methods such as METIS of k-means are not able to perform agglomeration of meshes describing heterogeneous domains and/or containing embedded interfaces. The unique possibility of performing this task with such methods is to do the agglomeration separately on the different regions and then merge them. However, the merging process of separate meshes is complex and time-consuming. Our model automates the procedure by exploiting the abilities of GNNs. Moreover, concerning domains with micro-structures, there is no chance of agglomerating them with state-of-the-art methods.
The GNN model proposed for the heterogeneous case is an extension of the GNN-AGGL-Enhanced model previously discussed. The most relevant change in the GNN architecture is the addition of a new input feature representing the physical parameter $\rho_{i}$ of the region to which the $i$-th element belongs. In this way we obtain an input feature matrix $X\in\mathbb{R}^{Nx5}$:
@@
where $(x_{i},y_{i},z_{i})$ are the $(x,y,z)$-coordinates of the $i$-th element’s center, while the $V_{i}$ is the element’s volume. As explained for the other quantities, the parameter $\rho_{i}$ would be re-scaled to obtain values between 00 and $1$. This practice reduces the variability of the data, making it easier for the network to learn the relative importance of this parameter on the final result.
It is well known that neural networks often struggle to handle sharp changes in input data that may lead to gradient instability and hinder the model’s generalization ability. Adding a new pre-processing operation can easily overcome this obstacle for heterogeneous domains. For this reason, we substitute in the normalization layer the last input feature with the average of such values in the adjacent elements. This step allows for avoiding sharp changes at the boundaries of distinct physical regions by smoothing out this transition.
The output of the GNN remains unchanged. It is $Y\in\mathbb{R}^{Nx2}$, where the first column represents the probability of belonging to $S_{1}$, while the second is the probability of being in $S_{2}$.
The idea behind the agglomeration algorithm is similar to that described in Algorithm 1. However, we introduce some modifications to better deal with the specific task of this model. Specifically, we insert an additional check before adding the agglomerated element to the final mesh to ensure the polytopal element connectivity. If it is not connected (an efficient algorithm for finding the connected components of a graph is reported in [50]) we perform an extra agglomeration step. The need to introduce this check is due to the network’s tendency to keep elements with the same physical parameter together, also in cases when they are not connected.
@@
Since this work is intended to be the first approach to GNN methods for heterogeneous mesh agglomeration, as a starting point we use a training dataset composed only of heterogeneous domains with a small number of parameters and regions. However, it is important to emphasize that this choice allows us to perform agglomeration also in the most general situations of computational domains with $S$ parameters or domains with micro-structures, as shown in the forthcoming numerical tests. Moreover, this choice allows us to evaluate the generalization capabilities of the model.
The training dataset contains $200$ tetrahedral meshes ($100$ with a single physical parameter and $100$ with two heterogeneous parameters). The validation dataset contains $50$ tetrahedral meshes ($25$ with a single physical parameter and $25$ with two different parameter regions). The choice of the number of elements in the dataset preserves the $80/20$ split between training and validation sets. The elements of the meshes in the two considered datasets are approximately $200$ for the coarsest mesh and 17 0001700017\,00017 000 for the finest one.
The training of the models has been performed in an unsupervised learning setting. We slightly modify the loss function by adding a further term to consider the physical feature. Hence, the new loss function includes a term involving the physical parameters. Consider the matrix $P\in\mathbb{R}^{N\times 2}$, that contains in the first column at the $i$-th row the ”physical” parameter $p_{i}$ and in the second one its complementary $1-p_{i}$ associated with the $i$-th mesh element. Then, the physical penalty part of the loss function reads:
@@
where $\odot$ denotes the element-wise multiplication and the summation is over all the entries of the resulting matrix. Finally, the loss function of the GNN is composed of the sum of the terms in Equations (2.1) and (1):
~$\tilde{\mathscr{L}}(W,P,Y,G)=\alpha\mathscr{P}(P,Y)+\beta\left(\mathscr{L}(W,Y,G)+\lambda\|W\|_{2}^{2}\right).$ (2)
The proposed loss function contains two parameters, $\alpha$ and $\beta$, to correctly balance the relevance of the normalized cut term and the one concerning the physical parameters. A hyperparameters tuning procedure has been used to balance the two terms. Indeed, by reducing the importance of the normalized cut, the model separates the physical classes correctly but we obtain low-quality agglomerated elements. Instead, by reducing the significance of the physical parameter term, we do not take into proper account the physical classes. The optimal values we found for that model are $\alpha=1.28$ and $\beta=0.00022$.
The training has been performed by adopting $150$ epochs of the Adam optimizer with a batch size of $4$. The learning rate is set to be $\gamma=10^{-4}$, and the weight decay parameter is $\lambda=10^{-5}$ and batch size $4$. In Figure 5 we report the decay of the loss function $\tilde{\mathscr{L}}$ evaluated on training and validation datasets.
0020202020404040406060606080808080100100100100120120120120140140140140001111222233334444Epochs~$\tilde{\mathscr{L}}(W,Y,G)$Loss function for the GNN-AGGL-Enhanced model on heterogeneous domainsTraining lossValidation loss
@@ 3 Numerical results: mesh agglomeration for homogeneous domains
In this section, after the introduction of some quality metrics for polyhedral meshes, we introduce a test dataset to compare the GNN model performance with the state-of-the-art methods (i.e. METIS and k-means). Moreover, we present an application to a mesh of human brain ventricles.
@@
    where $B(r)$ is a ball of radius $r$. We point out that inscribed spheres tangent to each of the polyhedron’s faces and circumscribed ones touching each of the polyhedron’s vertices might not exist for not convex polyhedra. For this reason, we instead use the largest sphere contained in the polyhedron as the inscribed one and the smallest sphere containing the polyhedron as the circumscribed one [51].
@@
The CR and UF quality metrics assume values between 00 and $1$. Moreover, if the elements are regular, these quality metrics are close to $1$. Finally, we introduce a metric to evaluate the performance of the agglomeration algorithm in terms of the difference between the volumes of the agglomerated mesh elements. VD assumes positive values and if the mesh is quasi-uniform, these values are close to 00.
@@ 3.1 Test case 1: agglomeration capabilities assessment of a test dataset
We construct a test dataset of $100$ tetrahedral meshes of different refinements, with $563$ elements for the coarsest mesh and 6 40464046\,4046 404 for the finest one. The dataset is composed as follows:
  1. 1.
    $30$ meshes of the unitary cube,
  2. 2.
    $70$ meshes of suitable portions of the unitary cube.
The computational domains are analogous to the ones used for training and we refer to Figure 3 that reports a mesh for each type. The results in the following sections are obtained using this test dataset. The considered agglomerated meshes have approximately $128$ elements, corresponding to $7$ bisection steps.
@@
Table 3: Test case 1: average values of the computed quality metrics (CR, UF, and VD) for the agglomerated grids reported in Figure 6, obtained with different agglomeration strategies (GNN-AGGL-Base, GNN-AGGL-Enhanced, METIS, k-Means).
@@
In Figure 6, we report a mesh from the test dataset together with the agglomerated meshes obtained using four different partitioning methods, namely GNN-AGGL-Base, GNN-AGGL-Enhanced, METIS, and k-means. We plot the agglomerated meshes with an ”exploded” view for visualization purposes. This allows us to see how tetrahedra are agglomerated together into the new mesh elements. We can see that GNN-AGGL-Enhanced and k-means grids look quite similar and the agglomerated elements seem to be shape regular. On the contrary, METIS and GNN-AGGL-Base ones are of lower quality, due to the presence of stretched elements.
The following Table 3 collects the average values over the full test set of the quality metrics (CR, UF, and VD) for the agglomerated grids obtained employing different agglomeration strategies (GNN-AGGL-Base, GNN-AGGL-Enhanced, METIS, and k-Means). Moreover, in Figure 7, we report the violin plots relative to the distributions of the quality metrics, for the considered agglomerated mesh. Looking at the results, it is clear that the GNN-AGGL-Base model is not able to create an agglomerated mesh of sufficiently good quality. As expected, due to the much greater variability in the three-dimensional contest, the performance of this model is significantly lower than the one obtained in the 2D case [27]. For this reason, we focus only on the GNN-AGGL-Enhanced model in the following test cases.
On the contrary, we can see that the GNN-AGGL-Enhanced model performs better than METIS. This can be easily explained because the latter considers only the information coming from the graph topology while GNN-AGGL-Enhanced takes in input some geometrical properties of the mesh. Observing also k-means, we can infer that the developed algorithm is the best in terms of UF, while it is slightly worse than the performance of k-means in terms of shape regularity (CR index). However, we observe that the CR index provides low values for all the considered methods.
Runtime Performance
@@
In Figure 8, we report the runtime performance of the proposed GNN-AGGL-Enhanced algorithm, compared with the state-of-the-art ones (METIS and k-Means). On the abscissa, we show the graph size, i.e. the number of meshes’ elements, and on the ordinate, we plot the time required for the graph’s bisection. The main advantage of our model with respect to the others is its non-iterative nature in performing graph bisection. The time benchmarking test set contains $15$ tetrahedral meshes of increasing size, from $24$ to 19 0491904919\ 04919 049 elements.
As expected, due to the above-mentioned advantage, the GNN-AGGL-Enhanced model outperforms both METIS and k-means in terms of runtime performance. Concerning the homogeneous domain agglomeration, the speed of the developed model is a key point in the discussion and allows us to consider the remarkable agglomeration capabilities (in terms of the final mesh’s shape regularity) of our model slightly worse than, for example, k-means.
@@ 3.2 Test case 2: agglomeration of a mesh of human brain ventricles
To further test the generalization capabilities of the developed algorithm, we apply it to a more complex domain than the unitary cubes and portions of them considered so far. In particular, we consider some human brain ventricles reconstructed starting from a Magnetic Resonance Imaging (MRI) scan taken from the OASIS database [52]. Brain ventricles are a network of interconnected cavities filled with cerebrospinal fluid. The domain is highly non-convex and presents many convoluted pathways, constrictions, and narrowed sections. However, due to the actual research on brain fluid dynamics, constructing a high-quality polyhedral mesh is of major interest [2, 53]. The starting mesh consists of 28 7062870628\,70628 706 tetrahedral elements, as shown in Figure 9.
@@
We agglomerated such a grid using GNN-AGGL-Enhanced, METIS, and k-means algorithms. The target mesh size of Algorithm 1 used for this test case is $h/10$, where $h$ is the diameter of the original mesh. In this way, we obtain three agglomerated meshes with approximately 300 elements. In Figure 9 we report the results of the agglomeration processes, also considering, a detail of the agglomerated mesh of the fourth ventricle and aqueduct of Sylvius.
@@
Table 4: Test case 2: average values of the computed quality metrics (CR and UF) for the agglomerated grids reported in Figure 9, obtained with different agglomeration strategies (GNN-AGGL-Enhanced, METIS, k-Means).
In Table 4 we report the average values of the quality metrics (CR and UF) computed on the agglomerated values. The VD metric is not reported because the starting mesh and as a consequence the agglomerated ones are not uniform due to the geometrical conformation of the domain. Due to this fact, the evaluation of the uniformity of the volumes of the elements is not significant for the test case. The three considered algorithms show similar results. This confirms that the proposed GNN-AGGL-Enhanced algorithm is as accurate as the state-of-the-art ones. Moreover, it indicates a good generalization capability of the GNN-AGGL-Enhanced model, considering that the computational domain is totally different from the ones included in the training set, both in terms of shape and dimensions.
@@ 4 Numerical results: mesh agglomeration for heterogeneous domains
In this section, after the introduction of an additional quality metric to analyze the agglomeration domains with heterogeneous physical parameters, we introduce some test datasets to evaluate the GNN-AGGL-Enhanced model. Moreover, we present two applications, the agglomeration of a domain containing microstructure and of a human brain.
To evaluate the shape regularity of the elements in the agglomerated meshes, we use the quality metrics introduced in Section 3. However, we need to introduce a new quantity, that will be later used to measure the ability of the algorithm to preserve the different physical regions separated in the agglomeration process. We define the Heterogeneous Elements (HE) as the percentage of elements in the agglomerated mesh that contain discontinuities in the physical parameters.
@@
The quantity should take values close to 00 as much as possible, to avoid the presence of agglomerated elements containing discontinuities in the physical parameters of the domain.

### v2

@@ 2.2 Architectures of GNN-agglomeration model for heterogeneous domains
In this section, we further extend the proposed approach to perform mesh agglomeration on heterogeneous domains with different physical parameters. Specifically, this model can be exploited in different settings such as:
  • Heterogeneous domains, where the main goal of the algorithm is to maintain as much as possible the distinction between the different material properties, still ensuring an acceptable shape for the agglomerated elements.
  • Domains with micro-structures, which is a more complex situation because there is no precise criterion to perform the agglomeration, and the algorithm aims to understand how to group the elements to obtain a better grid for numerical methods.
We remark that graph partitioning algorithms like METIS or k-means are not able to perform agglomeration automatically, respecting the underlying heterogeneities and/or the embedded interfaces/microstructures. The only possibility of performing this task with the above-mentioned methods is to do the agglomeration separately on the different regions and then merge them. However, the merging process of separate meshes is complex and time-consuming. Our model automates the procedure by exploiting the abilities of GNNs.
The GNN model proposed for the heterogeneous case is an extension of the GNN-AGGL-Enhanced model previously discussed. The most relevant change in the GNN architecture is the addition of a new input feature representing the physical parameter $\rho_{i}$ of the region to which the $i$-th element belongs. In this way we obtain an input feature matrix $X\in\mathbb{R}^{N\times 5}$:
@@
where $(x_{i},y_{i},z_{i})$ are the $(x,y,z)$-coordinates of the $i$-th element’s center, while the $V_{i}$ is the element’s volume. As explained for the other quantities, the parameter $\rho_{i}$ would be re-scaled to obtain values between 0 and $1$. This practice reduces the variability of the data, making it easier for the network to learn the relative importance of this parameter on the final result. It is well known that neural networks often struggle to handle sharp changes in input data that may lead to gradient instability and hinder the model’s generalization ability. Adding a new pre-processing operation can easily overcome this obstacle for heterogeneous domains. For this reason, we substitute in the normalization layer the last input feature with the average of such values in the adjacent elements. This step allows for avoiding sharp changes at the boundaries of distinct physical regions by smoothing out this transition. The output of the GNN remains unchanged. It is $Y\in\mathbb{R}^{N\times 2}$, where the first column represents the probability of belonging to $S_{1}$, while the second is the probability of being in $S_{2}$.
The idea behind the agglomeration algorithm is similar to that described in Algorithm 1. However, we introduce some modifications to better deal with the specific task of this model. Specifically, we add an additional check before adding the agglomerated element to the final mesh to ensure the polytopal element connectivity. If the resulting agglomerate is not connected (an efficient algorithm for finding the connected components of a graph is reported in [55]), we perform an extra agglomeration step. The need to introduce this check is due to the network’s tendency to keep elements with the same physical parameter together, even when they are not connected.
@@
Since this work is intended to be the first approach to GNN methods for heterogeneous mesh agglomeration, we use a training dataset composed only of heterogeneous domains with a small number of parameters and regions as a starting point. However, it is important to emphasize that this choice also allows us to perform agglomeration in the most general situations of computational domains with $S$ parameters or domains with micro-structures, as shown in the forthcoming numerical tests. Moreover, this choice allows us to evaluate the generalization capabilities of the model. The training dataset contains $200$ tetrahedral meshes ($100$ with a single physical parameter and $100$ with two heterogeneous parameters). The validation dataset contains $50$ tetrahedral meshes ($25$ with a single physical parameter and $25$ with two different parameter regions). The choice of the number of elements in the dataset preserves the $80/20$ split between training and validation sets. The elements of the meshes in the two considered datasets are approximately $200$ for the coarsest mesh and 17 00017\,000 for the finest one. The training of the models has been performed in an unsupervised learning setting. We slightly modify the loss function by adding a further term to consider the physical feature. Hence, the new loss function includes a term involving the physical parameters. Consider the matrix $P\in\mathbb{R}^{N\times 2}$, which contains in the first column at the $i$-th row the ”physical” parameter $p_{i}$ and in the second column its complement $1-p_{i}$ associated with the $i$-th mesh element. Then, the physical penalty part of the loss function reads:
@@
where $\odot$ denotes the element-wise multiplication and the summation is over all the entries of the resulting matrix. Finally, the loss function of the GNN is composed of the sum of the terms in Equations (2) and (3):
$\tilde{\mathscr{L}}(W,P,Y,G)=\alpha\mathscr{P}(P,Y)+\beta\left(\mathscr{L}(W,Y,G)+\lambda\|W\|_{2}^{2}\right).$ (4)
The proposed loss function contains two parameters, $\alpha$ and $\beta$, to correctly balance the relevance of the normalized cut term and the term related to physical parameters. A hyperparameter tuning procedure has been used to balance the two terms. Indeed, by reducing the normalized cut’s importance, the model separates the physical classes correctly, but we obtain low-quality agglomerated elements. Instead, by reducing the significance of the physical parameter term, we fail to account for the physical classes. The optimal values we found for that model are $\alpha=1.28$ and $\beta=2.2\times 10^{-4}$. The training has been performed by adopting $150$ epochs of the Adam optimizer with a batch size of $4$. The learning rate is set to be $\gamma=10^{-4}$, and the weight decay parameter is $\lambda=10^{-5}$ and batch size $4$. In Figure 5 we report the decay of the loss function $\tilde{\mathscr{L}}$ evaluated on training and validation datasets. The GNN training was carried out on a workstation equipped with an Intel Core i7-1270P CPU (12 cores, 16 threads, base clock 2.2GHz) and 32GB RAM. Concerning the training times, we have to consider a cost of $\simeq 500$ minutes for the complete training of the GNN-AGGL-Enhanced network. However, this cost is offline, so we need to perform it only once. This cost can be compensated with the lower online costs with respect to the METIS or k-means algorithms, as discussed in the next sections. Moreover, the training set can be limited, containing small meshes, due to the high generalization properties over larger meshes of the GNN architecture in partitioning [46]. This is a large advantage of the proposed method, which can be applied to agglomerate large meshes without an increase in the training costs.
[…]
@@ 3 Numerical results: mesh agglomeration for homogeneous domains
In this section, after the introduction of some quality metrics for polyhedral meshes, we introduce a test dataset to compare the performance of the GNN model with graph-partitioning methods like METIS and k-means. Moreover, we present an application to a mesh of human brain ventricles.
@@
    where $B(r)$ is a ball of radius $r$. We point out that inscribed spheres tangent to each of the polyhedron’s faces and circumscribed ones touching each of the polyhedron’s vertices might not exist for non-convex polyhedra. For this reason, we instead use the largest sphere contained in the polyhedron as the inscribed one and the smallest sphere containing the polyhedron as the circumscribed one [56].
@@
The CR and UF quality metrics assume values between 0 and $1$. Moreover, if the elements are regular, these quality metrics are close to $1$. Finally, we introduce a metric to evaluate the performance of the agglomeration algorithm in terms of the difference between the volumes of the agglomerated mesh elements. VD assumes positive values, and if the mesh is quasi-uniform, these values are close to 0.
@@ 3.1 Test case 1: agglomeration capabilities assessment of a test dataset
We construct a test dataset of $100$ tetrahedral meshes with different levels of refinements, ranging from $563$ elements for the coarsest mesh and 6 4046\,404 for the finest one. The dataset is composed as follows: 1) $30$ meshes of the unitary cube; 2) $70$ meshes of suitable portions of the unitary cube. The computational domains are analogous to the ones used for training, and we refer to Figure 3 that reports a mesh for each type. The results in the following sections are obtained using this test dataset. The considered agglomerated meshes have approximately $128$ elements, corresponding to $7$ bisection steps.
@@
Table 3: Test case 1: average values of the computed quality metrics (CR, UF, and VD) for the agglomerated grids reported in Figure 6, obtained with different agglomeration strategies (GNN-AGGL-Base, GNN-AGGL-Enhanced, METIS, k-Means). The colors are scaled starting from the minimum value of the metric (red) to the maximum one (green).
@@
In Figure 6, we report a mesh from the test dataset together with the agglomerated meshes obtained using four different partitioning methods, namely GNN-AGGL-Base, GNN-AGGL-Enhanced, METIS, and k-means. We plot the agglomerated meshes with an ”exploded” view for visualization purposes. This allows us to see how tetrahedra are agglomerated together into the new mesh elements. We can see that GNN-AGGL-Enhanced and k-means grids look similar. On the contrary, METIS and GNN-AGGL-Base ones are of lower quality, due to the presence of stretched elements. Table 3 collects the average values over the full test set of the quality metrics (CR, UF, and VD) for the agglomerated grids obtained employing different agglomeration strategies (GNN-AGGL-Base, GNN-AGGL-Enhanced, METIS, and k-Means). Moreover, in Figure 7, we report the violin plots relative to the distributions of the quality metrics for the considered agglomerated mesh. Looking at the results, it is clear that the GNN-AGGL-Base model is not able to create an agglomerated mesh of sufficiently good quality. As expected, due to the much greater variability in the three-dimensional context, the performance of this model is significantly lower than that obtained in the two-dimensional case [27]. For this reason, we focus only on the GNN-AGGL-Enhanced model in the following test cases. On the contrary, we can see that the GNN-AGGL-Enhanced model performs better than METIS. This can be easily explained because the latter considers only the information coming from the graph topology, while GNN-AGGL-Enhanced takes in input some geometrical properties of the mesh. Observing also k-means, we can infer that the developed algorithm seems to be the best in terms of UF, while it is slightly worse than k-means in terms of shape regularity (CR index). However, we observe that the CR index provides low values for all the considered methods.
We conclude this section by presenting some results on runtime performance.
@@
In Figure 8, we report the runtime performance of the proposed GNN-AGGL-Enhanced algorithm, and compare it with the analogous quantities obtained with METIS and k-Means algorithms. On the abscissa, we show the graph size, i.e., the number of mesh elements, and on the ordinate, we plot the time required for the graph’s bisection. The main advantage of our model with respect to the others is its non-iterative nature in performing graph bisection. The time benchmarking test set contains $15$ tetrahedral meshes of increasing size, from $24$ to 19 04919\ 049 elements. Due to the advantage mentioned above, the GNN-AGGL-Enhanced model seems to outperform both METIS and k-means in terms of runtime performance.
@@ 3.2 Test case 2: agglomeration of a mesh of human brain ventricles
To further test the generalization capabilities of the developed algorithm, we apply it to a more complex domain. In particular, we consider human brain ventricles reconstructed starting from a Magnetic Resonance Imaging (MRI) scan taken from the OASIS database [57]. Brain ventricles are a network of interconnected cavities filled with cerebrospinal fluid. The domain is highly non-convex and presents many convoluted pathways, constrictions, and narrowed sections. However, due to the actual research on brain fluid dynamics, constructing high-quality meshes is of major interest [2, 58]. The initial tetrahedral mesh consists of 28 70628\,706 tetrahedral elements, as shown in Figure 9.
@@
We agglomerated such a grid using GNN-AGGL-Enhanced, METIS, and k-means algorithms. The target mesh size of Algorithm 1 used for this test case is $h/10$, where $h$ is the diameter of the initial mesh. This choice leads to agglomerated meshes with approximately 300 elements. In Figure 9 we report the results of the agglomeration processes, also considering a detail of the agglomerated mesh of the fourth ventricle and aqueduct of Sylvius.
@@
Table 4: Test case 2: computed quality metrics (CR and UF) for the agglomerated grids reported in Figure 9, obtained with different agglomeration strategies (GNN-AGGL-Enhanced, METIS, k-Means). The colors are scaled starting from the minimum value of the metric (red) to the maximum one (green).
In Table 4, we report the quality metrics (CR and UF) computed on the agglomerated meshes. The VD metric is not reported because the starting mesh and, as a consequence, the agglomerated ones are not uniform due to the geometrical conformation of the domain. Due to this fact, the evaluation of the uniformity of the volumes of the elements is not significant for this test case. The three considered algorithms show similar results. This confirms that the proposed GNN-AGGL-Enhanced algorithm seems to be as accurate as METIS and k-means. Moreover, it indicates a good generalization capability of the GNN-AGGL-Enhanced model, considering that the computational domain is totally different from the ones included in the training set, both in terms of shape and dimensions.
@@ 4 Numerical results: mesh agglomeration for heterogeneous domains
In this section, after introducing an additional quality metric to test the proposed agglomeration algorithm in domains featuring heterogeneous physical parameters, we introduce some test datasets to evaluate the GNN-AGGL-Enhanced model. Moreover, we present as an application, the agglomeration of a mesh representing a domain with microstructure.
To evaluate the regularity of the agglomerated elements, we use the quality metrics introduced in Section 3. However, we need to introduce a new quantity that will be later used to measure the ability of the algorithm to preserve the different physical regions separated in the agglomeration process. We define the Heterogeneous Elements (HE) as the percentage of elements in the agglomerated mesh that contain discontinuities in the physical parameters.
@@
The quantity should take values close to 0 as much as possible, to avoid the presence of agglomerated elements containing discontinuities in the physical parameters.